OpenStudy (amonoconnor):

I'm confused as to how my textbook got the answer for the following problem. Can someone walk me through how they get the answer? I'm so close, but my answer isn't quite there and I'm not sure on the reasoning for how to get to the book's answer. Any and all help is greatly appreciated!

OpenStudy (anonymous):

why didnt you write it?

OpenStudy (amonoconnor):

Because it's unpleasant algebraic terms... just hold on. I'm putting it in the equation writer.

OpenStudy (nermaljean99):

So whats your problem exactly and what grade?

OpenStudy (amonoconnor):

The problem: \[\int\limits_{}^{}f^t(\sin(5^t))dt\] \[u = 5t\] \[du = 5^t[\log(5)]dt\] \[\frac{1}{\log(5)}*du = 5^tdt\]

OpenStudy (nermaljean99):

Blargh. ok hang on... lemme get my handy graph calculator out

OpenStudy (nermaljean99):

What grade is this stuff?!

OpenStudy (amonoconnor):

So... \[\int\limits_{}^{}5^t[\sin(5^t)]dt = \int\limits_{}^{}\sin(u)*\frac{1}{\log(5)}du\] \[= (-\cos(u))*\frac{u}{\log(5)}\] \[= –\cos(5^t)*\frac{5^t}{\log(5)} = \frac{–5tcos(5^t)}{\log(5)}\] That's what I got :/ But my book says it's wrong, and the only difference is that they don't have the coefficient of "5^t" in the numerator at the end, yet wolframalpha says that the antiderivative of 1/log(x) = x/log(x)... is "u" in this case not x, and wouldn't it be on the top at the end?

OpenStudy (amonoconnor):

This is college Calc 1:)

OpenStudy (nermaljean99):

Im still in high school geometry but ill try to help

OpenStudy (anonymous):

|dw

OpenStudy (amonoconnor):

@freckles @ganeshie8 Do you guys have a moment? :/

OpenStudy (nermaljean99):

are you trying to determine if its convergent? because it is not..

OpenStudy (amonoconnor):

No, I'm just evaluating the Integral:)

OpenStudy (nermaljean99):

Oh duh. Sorry I am a skimmer when I read XD I like to get straight to the work... ummm um um... ok hang on. let me go back and rework it then

OpenStudy (nermaljean99):

I am sorry hun, I dont understand it much myself

OpenStudy (nermaljean99):

Wish I could help

OpenStudy (amonoconnor):

No worries:) Anyone else?

OpenStudy (pawanyadav):

Integration of sin(u)/log5 du =-cos(u)/log5 +c

OpenStudy (pawanyadav):

You multiply it by u in numerator that was the mistake

OpenStudy (amonoconnor):

But why don't I consider the "unseen u^0" that's multiplied by 1 in the numerator, when integrating?

OpenStudy (amonoconnor):

Is this just a formula/rule thing to memorize?

OpenStudy (pawanyadav):

That' have no effect on the solution

OpenStudy (pawanyadav):

u^0=1. which don't effect anything

OpenStudy (amonoconnor):

Is the following not true, unless there's an identity or rule about the integration of logs? \[\int\limits_{}^{}\sin(u)*\frac{1}{\log(5)} = \int\limits_{}^{}\sin(u)*\frac{u^0}{\log(5)} = [(-\cos(u))*\frac{u^1}{\log(5)}]\]

OpenStudy (amonoconnor):

Exactly, u^0 = 1, meaning that when integrated, [u^(0+1)]/(0+1) / u^1 = u... ? Am I wrong?

OpenStudy (pawanyadav):

This is done when you have only Integration of u^0du or simply du

OpenStudy (pawanyadav):

In this question it's meaningless

OpenStudy (amonoconnor):

So it's a rule having to do with logs, and the antiderivatives of them?

OpenStudy (pawanyadav):

Any problem now

OpenStudy (pawanyadav):

I think you are clear now

OpenStudy (amonoconnor):

I think... thank you for your help!

OpenStudy (pawanyadav):

You are welcome

OpenStudy (amonoconnor):

MY 4th comment up is the summation of my question:)

OpenStudy (freckles):

\[\int\limits 5^t \sin(5^t) dt \\ \text{ Let } u=5^{t} \\ \text{ then } \frac{du}{dt}=5^t \ln(5) \\ du=5^t \ln(5) dt \\ \frac{1}{\ln(5)} du=5^t dt \\ \int\limits \sin(5^t) \cdot 5^t dt =\int\limits \sin(u) \frac{1}{\ln(5)} du \\ \text{ recall } \frac{1}{\ln(5)} \text{ is a constant } \\ \text{ and this constant is a constant multiple of that one function } \sin(u) \\ \\ \text{ since } \frac{1}{\ln(5)} \text{ is a constant multiple we can bring it outside } \\ \frac{1}{\ln(5)} \int\limits \sin(u) du\]

OpenStudy (freckles):

also... pretend we have \[\int\limits u du \text{ we say this is equal \to } \frac{u^2}{2}+C \\ \text{ \not } \int\limits u \cdot 1 du=\frac{u^2}{2} \cdot u+C \text{ however} \\ \text{ to integrate a product you could apply integration by parts} \\ \text{ while that is totally not needed for this example } \\ \text{ I'm just showing a possible way to integrate a product } \\ \int\limits u \cdot 1 du=\frac{u^2}{2} \cdot 1 -\int\limits \frac{u^2}{2} \cdot u +C \\ \int\limits u \cdot 1 du=\frac{u^2}{2}-\int\limits \frac{u^3}{2}+C\] so while you could put a one constant muliple in here it is really not helpful in this case I can show you a case where it is helpful for example \[\int\limits \ln(x) dx \\ \int\limits 1 \cdot \ln(x) dx \\ x \ln(x)-\int\limits x \frac{1}{x} dx \\ x \ln(x)-\int\limits 1 dx \\ x \ln(x)-x +C\] but anyways in general what I'm trying to get across from my examples is: \[\int\limits f(x) dx =\int\limits 1 \cdot f(x) dx \neq \int\limits 1 dx \cdot \int\limits f(x) dx\]

OpenStudy (freckles):

which is what you did

OpenStudy (freckles):

the slash through the equal sign means not equal if you haven't seen that before

OpenStudy (freckles):

oops one sec there is no mistake above but I meant to do that one integral the other way... \[\int\limits u \cdot 1 du=u \cdot u -\int\limits 1 \cdot u du \\ \int\limits u \cdot 1 du=u^2-\int\limits u du\] so actually it was helpful in that case I just chose the wrong one to integrate

OpenStudy (freckles):

\[\int\limits u du = u^2 - \int\limits u du +C \\ 2 \int\limits u du=u^2+C \\ \int\limits u du =\frac{1}{2} u^2 +\frac{1}{2} C \\ \int\limits u du =\frac{1}{2} u^2+K\]

OpenStudy (freckles):

I don't know i you have heard of integration by parts or not

OpenStudy (freckles):

\[\int\limits f(x) \cdot g'(x) dx=f(x) \cdot g(x) -\int\limits f'(x) \cdot g(x) dx\]

OpenStudy (amonoconnor):

Ah... this makes sense. Yes, I've heard of integration by parts, but haven't had a lot of practice yet. Thank you so much @freckles <3 ;D You rock!